# The early universe is ACT-ing warm

### v1

@@ II It is time to ACT now
P-ACT-BL has solidified the trend of pushing $n_{s}$ to higher values while combining the BK-18 results on a tiny upper bound on the tensor-to-scalar ratio. We would have to wait until the next data release of South Pole Telescope (SPT) to see if this trend of increasing $n_{s}$ while other missions will further constrain the upper bound on the amplitude of primordial gravitational waves. Our main message is that one does need to keep on fine-tuning existing cold inflation models by further adding parameters to account for future observations. Instead a principled approach would be to see what would happen if we were to add the dissipative corrections coming from WI and see what they predict the typical values for $n_{s}$ and $r$ are.
@@
The warm inflation model building program has focused on models built with the types of fields and symmetries that are as close as possible to those already found in the Standard Model, and have thus been experimentally observed. Developing such a goal has been possible due to two key steps. The first, we made in the early years of development of WI was the recognition that when the dissipative coefficient $\Upsilon$ is larger than the Hubble scale $H$, i.e. $\Upsilon>3H$, it is possible in this strong dissipative regime for the inflaton mass to be larger than the Hubble scale, $m_{\phi}>H$, and the inflaton field amplitude to remain below the Planck scale $\phi<m_{pl}$ Berera:1999ws ; Berera:2003yyp . Such a regime is then protected from quantum gravity corrections arising in higher dimensional operators and free of uncontrollable infrared problems. This is the minimal regime necessary to be consistent with the quantum field theory understood by the Standard Model. The second key step was recognizing, after much research by us, that for WI it is not Supersymmetry that best protects the inflaton mass from radiative and thermal corrections but rather the symmetries associated with the Nambu-Goldstone boson, leading to the first WI model based on this symmetry, the Warm Little inflaton model Bastero-Gil:2016qru . These two steps created a new path for inflation model building and the one followed in most WI model building today (for a more detailed discussion please see Berera:2023liv ).
II.1 The models
The first model we consider is the warm little inflaton model introduced in Ref. Bastero-Gil:2016qru , with the inflaton having direct interactions with fermion fields (we are going to refer to this model as WLIF for short) and with a quartic inflaton potential and interaction Lagrangian density given by
$\mathcal{L}_{\rm int}$ $=$ $-{g\over\sqrt{2}}(\phi_{1}+\phi_{2})\bar{\psi}_{1L}\psi_{1R}+i{g\over\sqrt{2}}(\phi_{1}-\phi_{2})\bar{\psi}_{2L}\psi_{2R}$ (10)
-- $h\sigma\sum_{i=1,2}\left(\bar{\psi}_{iL}\psi_{\sigma R}+\bar{\psi}_{\sigma L}\psi_{iR}\right),$
@@ II.1 The models
$\phi_{1}$ and $\phi_{2}$ are two complex scalar fields, with vacuum expectation values, $\langle\phi_{1}\rangle=\langle\phi_{2}\rangle\equiv M/\sqrt{2}$ and that are charged under the same $U(1)$ symmetry. The common phase is a pseudo-Goldstone boson that can be taken as the inflaton field. The remaining field content in Eq. (10), $\sigma$ is a scalar singlet, $\psi_{\sigma R}$ are chiral fermions carrying the same charge of $\psi_{1}$ and $\psi_{2}$, and $\psi_{\sigma L}$ has zero charge. The model satisfies the symmetries $\phi_{1}\leftrightarrow i\phi_{2}$ and $\psi_{1L,R}\leftrightarrow\psi_{2L,R}$, which protects the inflaton potential to receive any large quantum or thermal corrections. The dissipation coefficient evaluated in this case is
$\Upsilon=C_{T}T~{},\qquad C_{T}\simeq\alpha(h)g^{2}/h^{2}~{},$ (11)
where $\alpha(h)\simeq 3/[1-0.34\ln(h)]$. In the WLIF model, the form of the dissipation coefficient leads to a growing function $G(Q)$ in the power spectrum, which in general drives the tilt faster to a blue value ($n_{s}>1$). The simplest inflaton potential that can be considered in this case is a quartic monomial potential,
$V(\phi)=\frac{\lambda}{4}\phi^{4},$ (12)
which we will consider when presenting the results for this model.
The second model we analyze is a variant version of the WLIF model, with interaction Lagrangian density given by Bastero-Gil:2019gao
$\mathcal{L}_{\rm int}$ $=$ $-{1\over 2}g^{2}|\phi_{1}+\phi_{2}|^{2}|\chi_{1}|^{2}-{1\over 2}g^{2}|\phi_{1}-\phi_{2}|^{2}|\chi_{2}|^{2}$
$+$ $\!\!\!\!\!\sum_{\begin{subarray}{c}i\neq j=1,2\end{subarray}}\!\!\!\left(h\chi_{i}\bar{\psi}_{L}\psi_{R}\!+\!\mathrm{h.c.}\!-\!{\lambda\over 2}|\chi_{i}|^{4}\!-\!{\lambda^{\prime}}|\chi_{i}|^{2}|\chi_{j}|^{2}\!\right)\!,$
with principal scalar interactions instead of the femrion ones of the first model. We will denote this model by WLIS for short. As in the first model, the symmetries prevent the inflaton potential again from receiving large corrections. The dissipation coefficient derived in this case is found to be given by In this scalar field variant of the WLI model, the dissipation coefficient now becomes Bastero-Gil:2019gao
@@
$\Upsilon\simeq{4g^{4}\over h^{2}}{M^{2}T^{2}\over\tilde{m}_{\chi}^{3}}\left[1+{1\over\sqrt{2\pi}}\left({\tilde{m}_{\chi}\over T}\right)^{3/2}\right]e^{-\tilde{m}_{\chi}/T},$ (14)
where $\tilde{m}_{\chi}^{2}\simeq g^{2}M^{2}/2+\alpha^{2}T^{2}$, with $\alpha^{2}\simeq\left[h^{2}+2\lambda+\lambda^{\prime}\right]/12$. In the WLIS model, the dissipation coefficient behaves with a negative power in the temperature, $\Upsilon\propto 1/T$ for appropriate choices of parameters of the model. In this case, the function $G(Q)$ behaves like a decreasing function with the increase of $Q$. The simplest potential that can be studied in this case is just a quadratic monomial potential,
$V(\phi)=\frac{m_{\phi}^{2}}{2}\phi^{2},$ (15)
which is the choice of potential we consider here for the WLIS model.
@@
In the next section we present our results for these two models and contrast them with the recent P-ACT-LB-BK18 data.
II.2 Results
All of the results we are going to present are obtained by solving the WI equations given in the previous section, with the appropriate forms for the function $G(Q)$ correcting the power spectrum Eq. (6) which are derived for each of models considered (see, e.g. Refs. Graham:2009bf ; Bastero-Gil:2011rva for details and also Ref. Montefalcone:2023pvh for a public code to numerically determine $G(Q)$ for different forms of dissipation coefficients in WI). We also study the two limiting cases of considering the inflaton perturbations with negligible thermalization or fully thermalized, taking $n_{*}=0$ or $n_{*}=n_{BE}$ in Eq. (6). Finally, by setting the pivot scale, the point (given by the number of e-folds $N_{*}$) during inflation where it crosses the Hubble radius is determined explicitly by Das:2020xmh
$\frac{k_{*}}{a_{0}H_{0}}=e^{-N_{*}}\left[\frac{43}{11g_{s}(T_{\rm end})}\right]^{1/3}\frac{T_{0}}{T_{\rm end}}\frac{H_{*}}{H_{0}}\frac{a_{\rm end}}{a_{\rm reh}},$ (16)
[…]
Figure 1: The one-sigma and two-sigma constrains from Planck (orange shaded regions) and the combined P-ACT-LB results (violet shaded region) for $r$ and $n_{s}$, together with the results for the WLIF (panel a) and WLIS (panel b) models. The curves are the results obtained at different values for the dissipation ratio $Q$ at Hubble crossing (marked by the dots shown in the figure). The parameters considered in each case are the same as those used in Bastero-Gil:2016qru (for the WLIF model) and Bastero-Gil:2019gao (for the WLIS model).
@@ II.2 Results
In Fig. 1, we show the results for the two models, the WLIF and WLIS, obtained for the spectral tilt $n_{s}$ and for the tensor-to-scalar ratio $r$ as the value of $Q_{*}$ (evaluated at $N_{*}$) is varied (which corresponds to changes in the parameters and couplings of the two models). The behavior seen in the curves for the WI models in the figures in $r\times n_{s}$ is a consequence of the relative importance of each term in the power spectrum Eq. (full-power and how $Q$ and $T/H$ evolve (see, e.g. Ref. Das:2020lut for a discussion about the dynamical evolution for $Q$ and $T/H$ in WI and for their effect on $n_{s}$, see Ref. Das:2022ubr ).
@@
In addition to finding a slight spectral tilt towards scale invariance compared to Planck combined with BICEP/Keck data BICEP:2021xfz , the recent ACT data tend to slight favor a positive running (though still compatible with a null value), $\alpha_{s}=0.0062\pm 0.0052$. It is interesting to see that WI can also easily accommodate such values. The physical reason behind this is that the negative values for $\alpha_{s}$ in cold models come arise from the fact that the inflationary trajectories deviate from quasi-dS at the end of inflation when the shortest wavelength modes exit the horizon. While for WI, the energy density of the inflaton decreases continuously leading to a smooth transition to radiation domination. In Fig. 2, we show the results for the two models, the WLIF and WLIS, for the running $\alpha_{s}$ in terms of spectral tilt $n_{s}$, once again, different points in the curves correspond to different values of $Q_{*}$. In each case, the solid line corresponds to the results when neglecting $n_{*}$ in the equation for the scalar spectrum in WI or taking $n_{*}$ as a Bose-Einstein distribution (dashed line). The results fall well inside the one-sigma region found by ACT ACT:2025tim . It also show that for large $Q$ (as $n_{s}$ increases), the running tends to move from negative (in the weak dissipation regime of WI) towards positive values (corresponding to the strong dissipative WI regime), but still remains very small.
III Conclusions

### v2

@@ 3 It is time to ACT now
P-ACT-BL has solidified the trend of pushing $n_{s}$ to higher values while combining the BK-18 results on a tiny upper bound on the tensor-to-scalar ratio. We would have to wait until the next data release of South Pole Telescope (SPT) to see if this trend of increasing $n_{s}$ while other missions will further constrain the upper bound on the amplitude of primordial gravitational waves. Our main message is that one does need to keep on fine-tuning existing cold inflation models by further adding parameters to account for future observations. Instead, a principled approach would be to see what would happen if we were to add the dissipative corrections coming from WI and see what they predict the typical values for $n_{s}$ and $r$ are.
@@
The warm inflation model building program has focused on models built with the types of fields and symmetries that are as close as possible to those already found in the Standard Model, and have thus been experimentally observed. Developing such a goal has been possible due to two key steps. The first, we made in the early years of the development of WI was the recognition that when the dissipative coefficient $\Upsilon$ is larger than the Hubble scale $H$, i.e. $\Upsilon>3H$, it is possible in this strong dissipative regime for the inflaton mass to be larger than the Hubble scale, $m_{\phi}>H$, and the inflaton field amplitude to remain below the Planck scale $\phi<M_{\rm Pl}$ [41, 54]. Such a regime is then protected from quantum gravity corrections arising in higher-dimensional operators and is free of uncontrollable infrared problems. This is the minimal regime necessary to be consistent with the quantum field theory understood by the Standard Model. The second key step was recognizing, after much research by some of us [44, 55], that for WI it is not only Supersymmetry that best protects the inflaton mass from radiative and thermal corrections, but rather there can be other symmetries, particularly those reminiscent of beyond the standard model Higgs phenomenology, leading to the first WI model based on this symmetry, the Warm Little inflaton (WLI) model [28]. These two steps created a new path for inflation model building and the one followed in most WI model construction today (for a more detailed discussion, see [56, 57]).
3.1 The warm little inflaton model
In all of our analysis, we will consider a generalization of the WLI model, which unifies the models first introduced in refs. [28, 30] (see also  [58, 59] for applications and other generalizations). The WLI model is characterized by an interaction Lagrangian density given by
$\mathcal{L}_{\rm int}$ $=$ $-{g_{\phi\psi}\over\sqrt{2}}(\phi_{1}+\phi_{2})\bar{\psi}_{1L}\psi_{1R}+i{g_{\phi\psi}\over\sqrt{2}}(\phi_{1}-\phi_{2})\bar{\psi}_{2L}\psi_{2R}$ (3.1)
- ${1\over 2}g_{\phi\chi}^{2}|\phi_{1}+\phi_{2}|^{2}|\chi_{1}|^{2}-{1\over 2}g_{\phi\chi}^{2}|\phi_{1}-\phi_{2}|^{2}|\chi_{2}|^{2}$
- $h_{\sigma\psi}\sigma\sum_{i=1,2}\left(\bar{\psi}_{iL}\psi_{\sigma R}+\bar{\psi}_{\sigma L}\psi_{iR}\right)$
$+$ $\sum_{\begin{subarray}{c}i\neq j=1,2\end{subarray}}\left(h_{\chi\psi}\chi_{i}^{\dagger}\bar{\psi}_{\sigma L}\psi_{\sigma R}+\mathrm{h.c.}\right),$
where $\phi_{1}$ and $\phi_{2}$ are two complex scalar fields, with a symmetry breaking potential,
$V(\phi_{1},\phi_{2})=\sum_{i=1,2}\frac{\lambda_{i}}{4}\left(|\phi_{i}|^{2}-\frac{M^{2}}{2}\right)^{2},$ (3.2)
@@ 3.1 The warm little inflaton model
and vacuum expectation values, $\langle\phi_{1}\rangle=\langle\phi_{2}\rangle\equiv M/\sqrt{2}$. Both complex scalar fields are charged under the same $U(1)$ symmetry, which is spontaneously broken through (3.2). The common phase $\phi$ of the two fields, $\phi_{1}=Me^{i\phi/M}/\sqrt{2}$ and $\phi_{2}=Me^{-i\phi/M}/\sqrt{2}$, is a gauge singlet boson that can be taken as the inflaton field [^2] The phase $\phi$ being a gauge singlet, one can assign an arbitrary potential for it. This is different from axion-like (pseudo-Goldstone boson) potentials that are constrained by a shift symmetry.. The remaining field content in eq. (3.1), $\sigma$ is a scalar singlet, $\psi_{\sigma R}$ are chiral fermions carrying the same charge of $\psi_{1}$ and $\psi_{2}$, while $\psi_{\sigma L}$ has zero charge and $\chi_{1}$ and $\chi_{2}$ are complex scalar fields. The model satisfies the combined $Z_{4}$ and mirror symmetries [^3] These symmetries are reminiscent of the ones also seen in the little Higgs and twin Higgs type of models [60, 61]. For other recent connections of WI with standard model interactions and beyond, see, for example refs. [62, 63, 64, 65]. $\phi_{1}\leftrightarrow i\phi_{2}$, $\psi_{1L,R}\leftrightarrow\psi_{2L,R}$ and $\chi_{1}\to\chi_{2}$, which protects the inflaton potential to receive any large quantum or thermal corrections (see, e.g. [28, 30] for details).
When the Yukawa couling $g_{\phi\psi}$ is dominant over the biquadratic $g_{\phi\chi}$ one, we denote the model as the WLI fermionic (WLIF) model, while in the opposite case, with the biquadratic interactions between the $\phi_{i}$ and $\chi_{i}$ fields dominating over the Yukawa one between $\phi_{i}$ and $\psi_{i}$, we denote the model as the WLI scalar (WLIS) model. The dissipation coefficient evaluated in the WLIF case, in the high-temperature regime $g_{\phi\psi}M/T\ll 1$, and found to be given by [28]
$\Upsilon_{WLIF}=\alpha(h_{\sigma\psi})\frac{g_{\phi\psi}^{2}}{h_{\sigma\psi}^{2}}\,T,\quad g_{\phi\chi}\to 0,$ (3.3)
where $\alpha(h_{\sigma\psi})\simeq 3/[1-0.34\ln(h_{\sigma\psi})]$. In the case of the WLIS model, the dissipation coefficient was found to be given by [30]
@@
$\Upsilon_{WLIS}\simeq{4g_{\phi\chi}^{4}\over h_{\chi\psi}^{2}}{M^{2}T^{2}\over\tilde{m}_{\chi}^{3}}\left[1+{1\over\sqrt{2\pi}}\left({\tilde{m}_{\chi}\over T}\right)^{3/2}\right]e^{-\tilde{m}_{\chi}/T},\quad g_{\phi\psi}\to 0,$ (3.4)
where $\tilde{m}_{\chi}^{2}\simeq g_{\phi\chi}^{2}M^{2}/2+\alpha^{2}T^{2}$, with $\alpha^{2}\simeq\left[h_{\chi\psi}^{2}+\lambda_{\chi_{i}\chi_{j}}\right]/12$, where $\lambda_{\chi_{i}\chi_{j}}$ indicates possible inter- and intra-interactions among the complex scalars $\chi_{i}$ and that are allowed under the symmetries of the model.
In the WLIF model, the form of the dissipation coefficient leads to a growing function $G(Q)$ in the power spectrum, which in general drives the tilt faster to a blue value ($n_{s}>1$). On the other hand, in the WLIS model, the dissipation coefficient behaves with a negative power in the temperature, $\Upsilon_{WLIS}\propto 1/T$ for appropriate choices of model parameters. In this case, the function $G(Q)$ behaves as a decreasing function with increasing $Q$. These two types of behavior that appear depending on the temperature dependence of the dissipation coefficient will be explicitly verified when we analyze the results from the full model with the interactions (3.1), with the dissipation coefficient given by
$\Upsilon=\Upsilon_{WLIF}+\Upsilon_{WLIS}.$ (3.5)
@@
In the next section, we present our results using eq, (3.5) and contrast them with the recent P-ACT-LB-BK18 data and also with the earlier Planck data.
4 Results
An analysis of the individual separate dissipation coefficients (3.3) and (3.4) was done earlier in refs. [28] and [30], respectively. However, a unified analysis of the combined effect of both is still missing. We intend to fill this lacuna here.
We present the results considering the total dissipation coefficient eq. (3.5) for the WLI model eq. (3.1). Combining eqs. (3.3) and (3.4), we can write
$\Upsilon$ $=$ $\frac{g_{\phi\psi}^{2}}{h_{\sigma\psi}^{2}}\left\{\frac{3}{1-0.34\ln(h_{\sigma\psi})}\,T\right$.
$+$ $\left.b\,\frac{\left(2g_{\phi\chi}M\,T\right)^{2}}{\tilde{m}_{\chi}^{3}}\left[1+{1\over\sqrt{2\pi}}\left({\tilde{m}_{\chi}\over T}\right)^{3/2}\right]e^{-\tilde{m}_{\chi}/T}\right\},$
where $b$ is the ratio of the coupling constants,
$b=\frac{g_{\phi\chi}^{2}}{g_{\phi\psi}^{2}}\frac{h_{\sigma\psi}^{2}}{h_{\chi\psi}^{2}},$ (4.2)
and parameterizes the relative magnitude between the dissipation coefficients (3.3) and (3.4).
As already emphasized in the previous two sections, the strong dissipation regime of WI, where $Q>1$, is of particular interest. Hence, this is a regime that we will seek here. It will also guide us in the selection of the many possible parameters choices that can be made in the model (3.1). For instance, in the WLIF model, which was first studied in [28], the form of the dissipation coefficient (3.3) leads to a growing function $G(Q)$ in the power spectrum (2.5) and which in general drives the spectral tilt faster to a blue value ($n_{s}>1$) when $Q$ to goes to values above one. The simplest inflaton potential that can be considered in this case and that leads to values for the tensor-to-scalar ratio $r$ and tilt $n_{s}$ consistent with observations is a quartic monomial potential. However, the consistency with the observations favors the weak dissipation regime of WI, $Q<1$. Now in the case of the WLIS model, the dissipation coefficient (3.4) behaves mostly of the background dynamics with a negative power in the temperature, $\Upsilon\propto 1/T$ for appropriate choices of the model parameters. In this case, the function $G(Q)$ can display a decreasing function with increasing $Q$. In this regime of parameters, we can more easily have a strong dissipation with consistent values for $r$ and $n_{s}$. Furthermore, we can simply work with a quadratic monomial potential for the inflaton,
$V(\phi)=\frac{m_{\phi}^{2}}{2}\phi^{2}.$ (4.3)
4.1 Updated results for the WLIF and WLIS models
Before analyzing the complete WLI model with the combined dissipation coefficient (LABEL:UpsWLIFS), it is useful to revise the two cases WLIF and WLIS separately. The purpose for this is two-fold. Firstly, because they represent limiting cases for the complete model and will act as sanity checks of our main result. Secondly, because the analysis performed in the earlier references made use of simplified semi-analytical expressions to derive the observable quantities like $r$ and $n_{s}$, hence lacking precision as required when comparing with the most state of the art cosmological data. Here, this comparison between the predictions from the theory and observations is facilitated with the use of the recently released WI2easy code [29]. WI2easy is a dedicated code for WI dynamics, enabling efficient computation of both background evolution and curvature perturbations that makes possible precision comparisons between WI predictions and observational constraints.
(a) WLIF
(b) WLIS
(c) WLIF
(d) WLIS
Figure 1: The one-sigma and two-sigma constraints from Planck (orange shaded regions) and the combined P-ACT-LB results (violet shaded region) for $n_{s}$ as a function of $Q_{*}$ (panels a, b) and for $n_{s}\times r$ (panels c, d). Panels a and c are the results for the WLIF models (with a quartic inflaton potential), while panels b and d are for the WLIS model (with a quadratic inflaton potential). The shaded region on the left of panels a and b indicates the region for which $r>0.036$.
In fig. 1, we show the results for the two models, the WLIF and WLIS, obtained for the spectral tilt $n_{s}$ and for the tensor-to-scalar ratio $r$ as the value of $Q_{*}$ is varied, with all quantities evaluated at the Hubble radius crossing point $N_{*}$ for the pivot scale (see, e.g. ref. [29] for details). Note that varying $Q_{*}$ corresponds to changes in the parameters and couplings of the two models. To enable comparison with the earlier references on these models, to obtain the results shown in fig. 1 the parameters considered in each case were chosen to be the same as those used in [28] (for the WLIF model) and [30] (for the WLIS model) and also with the choices of a quartic inflaton potential $V=\lambda\phi^{4}/4$ for the WLIF model and a quadratic potential, eq. (4.3), for the WLIS model. Note that the WLIF model does not lead to satisfactory results, when compared with observations, when a quadratic potential is used (see also the results for the combined dissipation shown below) [^4] See also ref. [66], where this case of a quadratic potential with a dissipation coefficient that has a linear dependence on the temperature was recently analyzed..
@@ 4.1 Updated results for the WLIF and WLIS models
The behavior seen in the curves for the WI models in fig. 1 in $r\times n_{s}$ is a consequence of the relative importance of each term in the power spectrum eq. (2.5) and how $Q$ and $T/H$ evolve (see, e.g. ref. [67] for a discussion of the dynamical evolution for $Q$ and $T/H$ in WI and for their effect on $n_{s}$, see ref. [50]).
@@
The previous references for the WLIF and WLIS models have not analyzed the behavior for the running $\alpha_{s}$ of the spectral tilt. Since the recent ACT data tend to slightly favor a positive running (though still compatible with a null value), it is also opportune to also make this analysis of how $\alpha_{s}$ behaves with $Q$ for the WLIF and WLIS models. This result is shown in fig. 2.
@@
It is interesting to see that WI can also easily accommodate positive values for $Q\gg 1$. The physical reason behind this is that the negative values for $\alpha_{s}$ in cold models arise from the fact that the inflationary trajectories deviate from quasi-dS at the end of inflation when the shortest wavelength modes exit the horizon. However, for WI the energy density of the inflaton continuously decreases, leading to a smooth transition to radiation domination (which in most cases is accompanied by an increase of $Q$). In Fig. 2, the solid line corresponds to the results when neglecting $n_{*}$ in the equation for the scalar spectrum in WI or taking $n_{*}$ as a Bose-Einstein distribution (dashed line). The results fall well within the one-sigma region found by ACT [22] for $\alpha_{s}$. It also shows that for large $Q$, the running tends to move from negative (in the weak dissipation regime of WI) towards positive values (corresponding to the strong dissipative WI regime), while it still remains very small.
4.2 Results for the full WLI model
Let us now turn to the presentation of the results for the full model (3.1), with dissipation coefficient given by eq. (LABEL:UpsWLIFS). Once again, we seek parameters that allow for the strong dissipative regime of WI, while also using the simplest inflaton potential eq. (4.3). As a representative set of parameters, we use those that already allowed strong dissipation in the case of the WLIS model analyzed in fig. 1(a). That is, with fixed value $g_{\phi\chi}M=2.6\times 10^{-5}M_{\rm Pl}$, coupling $\alpha^{2}=1/8$ and $h_{\sigma\psi}=2$. The remaining parameters, for example $m_{\phi}$, the individual values for the coupling $g_{\phi\chi}$ and mass scale $M$, along with the remaining coupling constants in (3.1) are all self-consistently determined. In particular, $m_{\phi}$ is evaluated by requiring that the scalar of curvature power spectrum at the pivot scale $k_{*}=0.05/Mpc$ has the CMB amplitude normalization [73] $A_{s}\simeq 2.105\times 10^{-9}$. The coupling constants and the mass parameter $M$ can be determined for each value of $Q$ by associating them with the chosen value for the ratio $b$ defined in eq. (4.2) and also with the overall prefactor $g_{\phi\psi}^{2}/h_{\sigma\psi}^{2}$ in eq. (LABEL:UpsWLIFS), which in the WI2easy code, it is associated with the $C_{\Upsilon}$ constant defined in there. We consider three representative cases for the ratio of coupling constants defined in eq. (4.2), b=1, 10, 20b=1,\,10,\,20.
Figure 3: The $G(Q)$ function generated through WI2easy [29] for a dissipation coefficient eq. (LABEL:UpsWLIFS) for b=1, 10, 20b=1,\,10,\,20, and for the potential (4.3). We present the result for both nonthermal $n_{*}=0$ and fully thermalized $n_{*}=n_{\rm BE}$. The inset plot zoom in around the region $10^{-4}\lesssim Q\lesssim 1$.
In fig. 3 we show the numerically generated results for $G(Q)$ through WI2easy and that is necessary in the expression for the power spectrum (2.5). We consider both the cases of including or not the thermal distribution term $n_{*}$ in eq. (2.6). In appendix A we give a justification of when this term might contribute in the power spectrum in the case of the WLI model. The results presented in fig. 3 show that there is no significant differences for $G(Q)$ for the present model when considering $n_{*}=0$ or $n_{*}=n_{\rm BE}$, except in the region $10^{-4}\lesssim Q\lesssim 1$. We note that the results for $G(Q)$ are only sensitive to the presence or not of $n_{*}$ in the weak dissipative regime. In obtaining the numerical results for $G(Q)$, we have also explicitly taken into account a stochastic noise term in the radiation perturbation equation that comes as a consequence of the conservation of the total stress energy tensor. This term was first discussed in [48] and it also affects the perturbations (through $G(Q)$) in the region $10^{-4}\lesssim Q\lesssim 1$. Earlier fittings that were proposed for $G(Q)$, e.g. in ref. [47] and also more recently in [49], have not taken into account the possibility of having $n_{*}=0$ or the presence of the above stochastic term in the perturbation equation for the radiation. Those fitting results for $G(Q)$ should then be used with care. In all of our results presented here, including the ones in figs. 1 and 2, we have explicitly included the stochastic term in the perturbation equation for the radiation (see also ref. [29] for details).
(a) $b=1$
(b) $b=10$
(c) $b=20$
(d) $b=1$
(e) $b=10$
(f) $b=20$
Figure 4: The same as in fig. 1, but showing the results for the WLI model, with a quadratic inflaton potential and for the cases of b=1, 10b=1,\,10 and $20$. Panels a, b and c show $n_{s}$ as a function of $Q_{*}$, while panels d, e and f show $n_{s}\times r$.
In fig. 4, we present the results for $n_{s}$ and $r$ obtained for the WLI model, with the dissipation coefficient given by eq. (LABEL:UpsWLIFS). We consider again the results when b=1, 10, 20b=1,\,10,\,20. Other cases with $b<1$ tend only to allow values of $r$ and $n_{s}$ compatible with the observations when the inflaton is thermalized, i.e., $n_{*}=n_{\rm BE}$ in eq. (2.6), but even so this only happens in the very weak dissipative regime of WI, $Q\ll 1$. As the value of $b$ is increased, the dissipation coefficient (3.4) starts to become more dominant throughout the dynamics, favoring the strong dissipative regime of WI, consistent with what is shown in fig. 1 for the WLIS model.
(a) $b=1$
(b) $b=10$
(c) $b=20$
Figure 5: The running $\alpha_{s}$ for the WLI model as a function of the dissipation ratio $Q_{*}$ at Hubble crossing and with the dissipation coefficient (LABEL:UpsWLIFS).
Finally, in fig. 5, we show the results for the running $\alpha_{s}$ for the WLI model and for the three cases of ratio of coupling constants $b$ considered above. We again see the trend observed for the separated cases of the WLIF and WLIS models, where the running tends to become positive for $Q\gtrsim 100$.
5 Conclusions
In this paper, we have built upon earlier results obtained for the warm little inflaton model first studied in ref. [28] when considering the leading dissipative effects coming from fermion interactions with the inflaton and then in ref. [30] for scalar interactions, by further generalizing the model. We have presented an analysis not done before for the combined use of fermionic and scalar interactions, in which case the dissipation coefficient in the model is extended to eq. (LABEL:UpsWLIFS). Moreover, our results were computed using the recently released code WI2easy [29], which provides the best precision analysis to date of background dynamics and cosmological perturbations in WI. Representative sets of parameters for the model have been considered, for which we have evaluated the relevant inflationary observable quantities, the tensor-to-scalar ratio $r$, the spectral tilt $n_{s}$ and the running of the tilt $\alpha_{s}$. We then compared the results for these with both the earlier Planck data and the more recent ACT data. As a consequence of the ACT data showing a preference for slightly larger values for $n_{s}$ compared to earlier Planck data, our results show that this tends to favor larger values for the dissipation coefficient ratio $Q$ in WI for the model studied here.